# Distance-Domain Degrees of Freedom in Near-Field Region

### v2

@@ I Introduction
The drive toward sixth-generation (6G) communication systems has spurred interest in exploiting large antenna apertures and higher frequency bands to meet increasing data throughput and spectral efficiency demands [1]. As the aperture size increases and the wavelength decreases, the near-field region expands and the spherical-wave nature of electromagnetic propagation becomes significant. Consequently, future wireless systems are more likely to operate in these near-field regimes [2], which has increased research interest in near-field communications.
@@
Unlike far-field scenarios, where beamforming is mainly used to steer energy along specific angular directions, the near-field enables beam focusing that leverages both angular and distance domains. This capability permits spatial multiplexing of users that share the same angular direction if they are located at different distances. Such an approach could be particularly useful in ultra-dense scenarios where many users are aligned in the same direction, a situation that challenges traditional far-field beamforming methods. Several studies have investigated spatial multiplexing based on the distance domain; for instance, [3] and [4] demonstrated through simulations that two collinear users can be spatially multiplexed in a LoS near-field channel. However, these works did not delve into a theoretical explanation of channel characteristics for the effectiveness of such multiplexing. To address this gap, [5] and [6] showed that the channel correlation between collinear users tends to decrease as their separation increases in the near-field region, so that once the correlation falls below a certain threshold, multiplexing becomes feasible. Based on the closed‐form channel‐correlation model in [5], Kosasih et al[7] introduced the concept of a finite beam depth for spatial multiplexing in the distance domain. In the conventional far‐field regime, the beam depth tends to diverge to infinity—resulting in very high correlation between collinear users and hence precluding distance‐domain multiplexing. By contrast, the near‐field region can yield a finite beam depth, which is equivalent to low spatial correlation between two collinear users, thereby enabling spatial multiplexing. This finite beam depth has been characterized analytically for a variety of aperture geometries, including rectangular and circular apertures [7], modular apertures [8, 9], and sparse arrays [10, 11], highlighting the potential for distance‐domain spatial multiplexing by using various types of apertures.
@@
In this work, we investigate the distance-domain DoF, which can provide many useful insights into the spatial multiplexing performance and channel estimation. For point-to-point (P2P) MIMO, the distance-domain DoF establish the maximum number of parallel non-interfering channels between Tx array and Rx array, where Rx array spans solely over the distance domain. In the context of multi-user MIMO (MU-MIMO), the distance-domain DoF can establish the maximum number of users that can be spatially multiplexed with minimal interference when they are at the same angle but different distance. In the context of channel estimation, the distance-domain DoF can represent the minimum number of samples/sweeping beams to recover the distance-dependent channel, assuming known user’s angle. As a result, establishing this limit sharpens our understanding of distance-domain beamforming and channel estimation. While recent studies [3, 4, 5, 6, 7] demonstrate the feasibility and benefits of distance-domain spatial multiplexing, they do not characterize the theoretical limit on the spatial DoF available along the distance dimension.
@@
To obtain an closed-form expression of the distance-domain DoF, we consider an equivalent point-to-point MIMO (P2P-MIMO) channel between a base station as a continuous-aperture transmit array of arbitrary shape and a continuous-aperture arbitrarily long receive array that spans solely over the distance domain. The continuous array can be arbitrarily long, capturing scenarios where users span a wide range of distances. Moreover, we allow the base-station aperture to have arbitrary shape, reflecting practical design flexibility.
@@ I-A Related Works
The DoF of near-field channels have been extensively investigated in point-to-point (P2P) MIMO settings; therefore, it is important to delineate how our problem relates to and differs from that literature. In general, most analyses consider geometries in which the receive (Rx) aperture spans primarily the angular domain of the transmit (Tx) aperture. Prior work has largely focused on configurations where one aperture covers the angular extent of the other, including linear arrays [12, 13, 14], rectangular arrays [15, 16, 17, 18, 19, 20], circular array [21] and more general apertures [22, 23, 24, 25].
In contrast to this extensive treatment of angular-domain DoF, the distance-domain DoF has received comparatively little attention. Some works do consider cases where the Rx aperture varies mainly along the distance domain of the Tx array [26, 27, 28, 29]. In [26, 27], the DoF between two rectangular apertures with mutually perpendicular orientations is analyzed. Although the Rx aperture in these studies spans the distance domain, it also extends over the angular domain. Hence, the DoF expression in [26, 27] is a combination of the distance-domain and angular-domain DoF, which prevents us from understanding the true values of the distance-domain DoF. In [28, Sec. IV-B], two linear arrays with perpendicular orientations are considered; this setup can capture distance-domain DoF for collinear users when the base station is modeled as a linear array, but it does not generalize to arbitrary Tx aperture shapes. Moreover, [29] derives a closed-form DoF expression for a rectangular Tx aperture and two collinear users at fixed positions. While insightful, the restriction to two predetermined users does not address the broader question studied here: determining the distance-domain DoF, equivalently maximum number of collinear users that can be spatially multiplexed solely via distance separation. Finally, all the works in [26, 28, 27, 29] assume Tx–Rx separations larger than the arrays dimensions, which does not reflect scenarios with collinear users distributed over a wide distance span.
@@ I-B1 System Model
  • Source (represents base station): The source is a Tx continuous-aperture array and may have arbitrary shape and may include gaps (capturing modular and sparse arrays). The only constraint is its inner and outer size, specified by the maximum radial extent from the array-center coordinate to its boundary.
  • Observation region (represents users): The observation region is a Rx continuous collinear array whose length can be arbitrarily large, which can represent multiple collinear users separated over a long distance span.
In contrast, most P2P-MIMO analyses target two specific, finite arrays (e.g., ULA/URA/UCA) and Rx size is smaller than the Tx–Rx separation, which cannot directly model the BS–multiuser channel with collinear, distance-separated users.
@@ I-B2 Technical Contributions
    where $p_{\min}$ and $p_{\max}$ denote the minimum and maximum Tx radial distances from the array center, and $r_{\min}$ and $r_{\max}$ denote the minimum and maximum Rx distances (along the collinear array) from the same center. The expression shows that the DoF depends only on the extreme edges of the apertures, not on the detailed Tx shape.
@@
The remainder of the paper is organized as follows. In Section II, we describe the system model, including the geometrical setup of Tx and Rx arrays, the near-field channel model. Section III presents the definition of distance-domain DoF and its implications in practice. Section IV presents the mathematical framework and derives a closed-form expression for the distance-domain DoF in the ideal case of Tx array and Rx array, with Rx array located in the broadside of Tx array  [^1] The conference version of this paper including Section II, III-A and IV is accepted for publication in IEEE Globecom 2025 [30].. Section V extends the analysis to the non-broadside configuration of the Rx array. Section VI extends the analysis to the modular Tx array. Finally, Section VII concludes the paper.
@@ II-A2 Rx linear array (can represent collinear users)
    where the distance span $r_{\max}-r_{\min}$ may be arbitrarily large and $r_{\max}-r_{\min}$ can possibly approach infinity. This contrasts with common assumptions under which $r_{\max}-r_{\min}\ll r_{\max}$, so the Tx–Rx separation is effectively a single distance [17, 16]. The infinite boundary is require to yield an upper bound on the distance-domain DoF for multi-user MIMO or channel estimation, where the distance span can be large. Moreover, $r_{\max}$ may approach the far field. Since the Fresnel approximation improves with increasing distance, our framework naturally covers joint near-field and far-field regimes.
@@ II-B Near-Field Channel Model
We consider a LoS channel between the Tx CAP array $\mathcal{P}$ and the Rx CAP linear array $\mathcal{Q}$. A full vector-wave description employs the free-space dyadic Green’s function [19]:
@@
where $k=2\pi/\lambda$, $h(\mathbf{q},\mathbf{p})$ is the scalar free-space Green’s function [19]
@@
When the minimum Tx–Rx distance is larger than the Tx-aperture size (often termed the Bjohnson distance [7]), each differential Tx element illuminates the Rx line with nearly uniform amplitude. Under this uniform-amplitude assumption and the Fresnel approximation, for $\mathbf{q}=r\,\mathbf{k}$ with $\|\mathbf{k}\|=1$,
@@ III-B Implications of distance-domain DoF for MU–MIMO and Channel Estimation
Figure 2: Implication of distance-domain DoF for users distributed within the distance span $r_{\min},r_{\max}]$
@@ III-B2 Minimum distance-domain sampling for near-field training
As a result, $\mathbf{h}(\mathbf{q})$ lies, up to the accuracy set by $\varepsilon$, in an $N_{\varepsilon}$-dimensional subspace. Consequently, any channel estimation technique that aims to approximate $\mathbf{h}(\mathbf{q})$ with a $C$-dimensional subspace by using the following codebook $\mathbf{W}\in\mathbb{C}^{M\times C}$, i.e., [31]
@@ IV-A Formulation of Integral Operator with Convolution Kernel
$\frac{1}{rr^{\prime}}\iint_{\mathscr{P}}\exp\!\Bigg(-\frac{j2\pi}{\lambda}\frac{\|\mathbf{p}\|^{2}-(\mathbf{k}^{T}\mathbf{p})^{2}}{2}\Big(\frac{1}{r}-\frac{1}{r^{\prime}}\Big)\Bigg)d\mathbf{p},$ (26)
@@
${v}(r,r^{\prime})$ $\approx\frac{1}{rr^{\prime}}\iint_{\mathscr{P}}\exp\!\Bigg(-\frac{j2\pi}{\lambda}\frac{\|\mathbf{p}\|^{2}}{2}\Big(\frac{1}{r}-\frac{1}{r^{\prime}}\Big)\Bigg)d\mathbf{p}.$ (27)
We cannot apply Landau’s Theorem [32] into the operator $\mathcal{V}$ since its kernel ${v}(r,r^{\prime})$ in (26) is not convolutional. This is because the amplitude ${1}/{rr^{\prime}}$ is not constant and $r$ can vary over a large distance. This is different from other P2P MIMO channel, where the element-wise distance between Tx and Rx arrays are relatively constant. To address this problem, a simple trick is to perform the following variable change:
@@
$g(t,t^{\prime})=g(t-t^{\prime})$ $=\iint_{\mathscr{P}}\exp\!\Big(-\frac{j2\pi}{\lambda}\frac{\|\mathbf{p}\|^{2}}{2}(t-t^{\prime})\Big)d\mathbf{p},$
@@
Lemma 1 indicates that the number of dominant eigenvalues of $\mathcal{V}$} can be deduced by analyzing $\mathcal{G}$. Moreover, because $g(t,t^{\prime})$ is a Hermitian convolution kernel, the number of dominant eigenvalues of $\mathcal{G}$ can be inferred from Landau’s Theorem (or Szego’s Theorem) [32]:

### v3

@@ I Introduction
The drive toward sixth-generation (6G) communication systems has spurred interest in exploiting large antenna apertures and higher frequency bands to meet increasing data throughput and spectral efficiency demands [liu2024near]. As the aperture size increases and the wavelength decreases, the near-field region expands and the spherical-wave nature of electromagnetic propagation becomes significant. Consequently, future wireless systems are more likely to operate in these near-field regimes [zhang2020prospective], which has increased research interest in near-field communications.
@@
Unlike far-field scenarios, where beamforming is mainly used to steer energy along specific angular directions, the near-field enables beam focusing that leverages both angular and distance domains. This capability permits spatial multiplexing of users that share the same angular direction if they are located at different distances. Such an approach could be particularly useful in ultra-dense scenarios where many users are aligned in the same direction, a situation that challenges traditional far-field beamforming methods. Several studies have investigated spatial multiplexing based on the distance domain; for instance, [zhang2022beam] and [li2023near] demonstrated through simulations that two collinear users can be spatially multiplexed in a LoS near-field channel. However, these works did not delve into a theoretical explanation of channel characteristics for the effectiveness of such multiplexing. To address this gap, [wu2023multiple] and [wu2023enabling] showed that the channel correlation between collinear users tends to decrease as their separation increases in the near-field region, so that once the correlation falls below a certain threshold, multiplexing becomes feasible. Based on the closed‐form channel‐correlation model in [wu2023multiple], Kosasih et al[kosasih2024finite] introduced the concept of a finite beam depth for spatial multiplexing in the distance domain. In the conventional far‐field regime, the beam depth tends to diverge to infinity—resulting in very high correlation between collinear users and hence precluding distance‐domain multiplexing. By contrast, the near‐field region can yield a finite beam depth, which is equivalent to low spatial correlation between two collinear users, thereby enabling spatial multiplexing. This finite beam depth has been characterized analytically for a variety of aperture geometries, including rectangular and circular apertures [kosasih2024finite], modular apertures [kosasih2024achieving, li2024multi], and sparse arrays [chen2024near, zhou2024sparse], highlighting the potential for distance‐domain spatial multiplexing by using various types of apertures.
@@
In this work, we investigate the distance-domain DoF, which can provide many useful insights into the spatial multiplexing performance and channel estimation. For point-to-point (P2P) MIMO, the distance-domain DoF establish the maximum number of parallel non-interfering channels between Tx array and Rx array, where Rx array spans solely over the distance domain. In the context of multi-user MIMO (MU-MIMO), the distance-domain DoF can establish the maximum number of users that can be spatially multiplexed with minimal interference when they are at the same angle but different distance. In the context of channel estimation, the distance-domain DoF can represent the minimum number of samples/sweeping beams to recover the distance-dependent channel, assuming known user’s angle. As a result, establishing this limit sharpens our understanding of distance-domain beamforming and channel estimation. While recent studies [zhang2022beam, li2023near, wu2023multiple, wu2023enabling, kosasih2024finite] demonstrate the feasibility and benefits of distance-domain spatial multiplexing, they do not characterize the theoretical limit on the spatial DoF available along the distance dimension.
@@
To obtain an closed-form expression of the distance-domain DoF, we consider an equivalent point-to-point MIMO (P2P-MIMO) channel between a base station as a continuous-aperture transmit array of arbitrary shape and a continuous-aperture arbitrarily long receive array that spans solely over the distance domain. The continuous array can be arbitrarily long, capturing scenarios where users span a wide range of distances. Moreover, we allow the base-station aperture to have arbitrary shape, reflecting practical design flexibility. We focus on the LoS component, which typically dominates in mmWave/THz bands due to significant NLoS attenuation [priebe2013stochastic]. Since multipath scattering typically increases channel rank by introducing virtual sources [tse2005fundamentals], our LoS analysis establishes a fundamental DoF lower bound characterizing performance in sparse-scattering environments.
@@ I-A Related Works
The DoF of near-field channels have been extensively investigated in point-to-point (P2P) MIMO settings; therefore, it is important to delineate how our problem relates to and differs from that literature. In general, most analyses consider geometries in which the receive (Rx) aperture spans primarily the angular domain of the transmit (Tx) aperture. Prior work has largely focused on configurations where one aperture covers the angular extent of the other, including linear arrays [xie2023performance, pu2014effects, torkildson2011indoor], rectangular arrays [hu2018beyond, dardari2020communicating, sun2021small, pizzo2022landau, do2023parabolic, gong2024holographic, wang2025analytical, guo2025impact], circular array [xu2017degrees] and more general apertures [yuan2021electromagnetic, shen2023electromagnetic, xu2023exploiting, miller2000communicating, pizzo2022nyquist, gustafsson2025degrees].
In contrast to this extensive treatment of angular-domain DoF, the distance-domain DoF has received comparatively little attention. Some works do consider cases where the Rx aperture varies mainly along the distance domain of the Tx array [decarli2021communication, ruiz2023degrees, ding2022degrees, chen2025degrees]. In [decarli2021communication, ruiz2023degrees], the DoF between two rectangular apertures with mutually perpendicular orientations is analyzed. Although the Rx aperture in these studies spans the distance domain, it also extends over the angular domain. Hence, the DoF expression in [decarli2021communication, ruiz2023degrees] is a combination of the distance-domain and angular-domain DoF, which prevents us from understanding the true values of the distance-domain DoF. In [ding2022degrees, Sec. IV-B], two linear arrays with perpendicular orientations are considered; this setup can capture distance-domain DoF for collinear users when the base station is modeled as a linear array, but it does not generalize to arbitrary Tx aperture shapes. Moreover, [chen2025degrees] derives a closed-form DoF expression for a rectangular Tx aperture and two collinear users at fixed positions. While insightful, the restriction to two predetermined users does not address the broader question studied here: determining the distance-domain DoF, equivalently maximum number of collinear users that can be spatially multiplexed solely via distance separation. Finally, all the works in [decarli2021communication, ding2022degrees, ruiz2023degrees, chen2025degrees] assume Tx–Rx separations larger than the arrays dimensions, which does not reflect scenarios with collinear users distributed over a wide distance span.
@@ I-B1 System Model
  • Source (represents base station): The source (Tx array) may have arbitrary shape and may include gaps (capturing modular and sparse arrays). The only constraint is its inner and outer size, specified by the maximum radial extent from the array-center coordinate to its boundary.
  • Observation region (represents users): The observation region (Rx array) is a collinear array whose length can be arbitrarily large, which can represent multiple collinear users separated over a long distance span.
  • Continuous model: We idealize both Tx and Rx as continuous-aperture (CAP) arrays. While being hardware-challenging, it serves as an information-theoretic benchmark: the CAP DoF establishes an upper bound on the DoF of any sampled (discrete) implementation. [^1] Extension of this work, i.e., the problem of sampling the continuous source and observation to maximize the distance-domain DoF, is under preparation for publication [duong2025sampling]. This abstraction also yields clean insight into multi-user MIMO and channel-estimation limits.
In contrast, the literature focuses on P2P-MIMO with two specific, finite arrays (e.g., ULA/URA/UCA) and Rx size is smaller than the Tx–Rx separation, which cannot model the distance-domain DoF with wide distance span.
@@ I-B2 Technical Contributions
    where $p_{\min}$ and $p_{\max}$ denote the minimum and maximum Tx radial distances from the array center, and $r_{\min}$ and $r_{\max}$ denote the minimum and maximum Rx distances (along the collinear array) from the same center. The expression shows that the DoF depends only on the extreme edges of the apertures, not on the detailed Tx shape and aperture’s area.
@@
I-B3 Distinction from Angular-Domain DoF Results
It is worth to distinguish the distance-domain DoF derived in (1) from the well-established angular-domain DoF results found in the literature (e.g., Pizzo et al. [pizzo2022landau] and Ruiz-Sicilia et al. [ruiz2023degrees]). Given the extensive literature on angular-domain limits, one might intuitively assume the our proposed distance-domain DoF is merely a special case of the angular-domain DoF, obtainable by adjusting orientation parameters in the standard formula [torkildson2011indoor, pizzo2022landau, miller2000communicating, xie2023performance, pu2014effects, ruiz2023degrees, do2023parabolic, xu2017degrees, yuan2021electromagnetic, xu2023exploiting, guo2025impact]:
$\mathrm{DoF}_{\text{angular}}\approx\frac{\alpha A_{t}A_{r}}{(\lambda D)^{n}}.$ (2)
where $\alpha$ is coefficient determined via orientation of Tx and Rx arrays, $A_{t}$ and $A_{r}$ are the physical areas of Tx and Rx arrays, $D$ is the Tx–Rx separation (treated as a single representative distance), $n$ depends on aperture dimensionality.
It can be demonstrated that (2) from [pizzo2022landau, ruiz2023degrees] cannot yield the distance-domain DoF for two fundamental reasons:
  1. 1.
    Dependence on Area vs. Radial Extent: The angular-domain DoF in (2) scales linearly with the aperture areas ($A_{t},A_{r}$). In stark contrast, our derived distance-domain DoF (1) depends on the difference of squares of the radial extents ($p_{\max}^{2}-p_{\min}^{2}$) and difference of inverse distance ($r^{-1}_{\min}-r^{-1}_{\max}$) and not on the aperture areas.
  2. 2.
    Distance Spanning: Angular-domain models typically assume the Rx array dimensions are small relative to the transmission distance, effectively treating the Tx-Rx separation $D$ as a constant. Our model captures the wide-span regime where the Rx array extends from $r_{\min}$ to $r_{\max}$. Attempting to apply (2) by letting the Rx area grow along the distance axis leads to unbounded DoF growth, whereas our result correctly approaches a finite saturation limit as $r_{\max}\to\infty$.
The remainder of the paper is organized as follows. In Section II, we describe the system model, including the geometrical setup of Tx and Rx arrays, the near-field channel model. Section III presents the definition of distance-domain DoF and its implications in practice. Section IV presents the mathematical framework and derives a closed-form expression for the distance-domain DoF in the ideal case of Tx array and Rx array, with Rx array located in the broadside of Tx array  [^2] Parts of this paper including Section II, III-A and IV was presented in IEEE Globecom 2025 [son2025].. Section V extends the analysis to the non-broadside configuration of the Rx array. Section VI extends the analysis to the modular Tx array. Finally, Section VII concludes the paper.
@@ II-A2 Rx linear array (can represent collinear users)
    where the distance span $r_{\max}-r_{\min}$ may be arbitrarily large and $r_{\max}-r_{\min}$ can possibly approach infinity. This contrasts with common assumptions under which $r_{\max}-r_{\min}\ll r_{\max}$, so the Tx–Rx separation is effectively a single distance [do2023parabolic, pizzo2022landau]. The infinite boundary is require to yield an upper bound on the distance-domain DoF for multi-user MIMO or channel estimation, where the distance span can be large. Moreover, $r_{\max}$ may approach the far field. Since the Fresnel approximation improves with increasing distance, our framework naturally covers joint near-field and far-field regimes.
@@ II-B Near-Field Channel Model
We consider a LoS channel between the Tx CAP array $\mathcal{P}$ and the Rx CAP linear array $\mathcal{Q}$. A full vector-wave description employs the free-space dyadic Green’s function [wang2025analytical]:
@@
where $k=2\pi/\lambda$, $h(\mathbf{q},\mathbf{p})$ is the scalar free-space Green’s function [wang2025analytical]
@@
When the minimum Tx–Rx distance is larger than the Tx-aperture size (often termed the Bjohnson distance [kosasih2024finite]), each differential Tx element illuminates the Rx line with nearly uniform amplitude. Under this uniform-amplitude assumption and the Fresnel approximation, for $\mathbf{q}=r\,\mathbf{k}$ with $\|\mathbf{k}\|=1$,
@@ III-B Implications of Distance-Domain DoF for MU–MIMO and Channel Estimation
Figure 2: Implication of distance-domain DoF for users distributed within the distance span $r_{\min},r_{\max}]$. The linear array (red) represents the region where single-antenna users can be located.
@@ III-B2 Minimum distance-domain sampling for near-field training
As a result, $\mathbf{h}(\mathbf{q})$ lies, up to the accuracy set by $\varepsilon$, in an $N_{\varepsilon}$-dimensional subspace. Consequently, any channel estimation technique that aims to approximate $\mathbf{h}(\mathbf{q})$ with a $C$-dimensional subspace by using the following codebook $\mathbf{W}\in\mathbb{C}^{M\times C}$, i.e., [cui2022channel]
@@ IV-A Formulation of Integral Operator with Convolution Kernel
$\frac{1}{rr^{\prime}}\iint_{\mathcal{P}}\exp\!\Bigg(-\frac{j2\pi}{\lambda}\frac{\|\mathbf{p}\|^{2}-(\mathbf{k}^{T}\mathbf{p})^{2}}{2}\Big(\frac{1}{r}-\frac{1}{r^{\prime}}\Big)\Bigg)d\mathbf{p},$ (27)
@@
${v}(r,r^{\prime})$ $\approx\frac{1}{rr^{\prime}}\iint_{\mathcal{P}}\exp\!\Bigg(-\frac{j2\pi}{\lambda}\frac{\|\mathbf{p}\|^{2}}{2}\Big(\frac{1}{r}-\frac{1}{r^{\prime}}\Big)\Bigg)d\mathbf{p}.$ (28)
We cannot apply Landau’s Theorem [landau1975szego] into the operator $\mathcal{V}$ since its kernel ${v}(r,r^{\prime})$ in (27) is not convolutional. This is because the amplitude ${1}/{rr^{\prime}}$ is not constant and $r$ can vary over a large distance. This is different from other P2P MIMO channel, where the element-wise distance between Tx and Rx arrays are relatively constant. To address this problem, a simple trick is to perform the following variable change:
@@
$g(t,t^{\prime})=g(t-t^{\prime})$ $=\iint_{\mathcal{P}}\exp\!\Big(-\frac{j2\pi}{\lambda}\frac{\|\mathbf{p}\|^{2}}{2}(t-t^{\prime})\Big)d\mathbf{p},$
@@
Lemma 1 indicates that the number of dominant eigenvalues of $\mathcal{V}$} can be deduced by analyzing $\mathcal{G}$. Moreover, because $g(t,t^{\prime})$ is a Hermitian convolution kernel, the number of dominant eigenvalues of $\mathcal{G}$ can be inferred from Landau’s Theorem (or Szego’s Theorem) [landau1975szego]: